OpenStudy (anonymous):

Find the derivative of the function. g(t) = e^(−9/t^(9)) g'(t) =

OpenStudy (bahrom7893):

g'(t) = e^(−9/t^(9)) * (−9/t^(9))'

OpenStudy (dumbcow):

chain rule :)

OpenStudy (anonymous):

chain rule :(

OpenStudy (bahrom7893):

(−9/t^(9))' = (−9*t^(-9))'=81t^(-8)

OpenStudy (anonymous):

what is the 81t^(-8)?

OpenStudy (anonymous):

it's not g'(t), right?

OpenStudy (anonymous):

zarkon, please help!

OpenStudy (zarkon):

g(t) = e^(−9/t^(9)) \[g(t)=e^{\frac{-9}{t^9}}=e^{-9t^{-9}}\] \[g'(t)=e^{-9t^{-9}}(-9)(-9)t^{-9-1}=e^{-9t^{-9}}81t^{-10}\]

OpenStudy (zarkon):

ok?

OpenStudy (anonymous):

where did the second (-9) come from?

OpenStudy (zarkon):

\[\frac{d}{dt}\left(t^{-9}\right)=-9t^{-10}\]